OpenStudy (anonymous):

Prove this identity sin^2 Θcsc^2 Θ = sin^2 Θ = cos^2 Θ

myininaya (myininaya):

Think you have too many equality symbols

OpenStudy (anonymous):

That is what the question I got for homework is asking. I don't understand it.

OpenStudy (mertsj):

You copied it wrong. It should be: \[\sin ^2x \csc ^2x-\sin^2x=\cos^2x\]

myininaya (myininaya):

\[\sin^2(\theta)+\cos^2(\theta)=1\] \[1+\cot^2(\theta)=\csc^2(x) => \csc^2(x)-1=\cot^2(x)\] So maybe it is suppose to say \[\sin^2(\theta) \csc^2(\theta)-\sin^2(\theta)=\cos^2(\theta)\] \[\sin^2(\theta)(\csc^2(\theta)-1)=\sin^2(\theta) \cdot \cot^2(x)=\sin^2(\theta) \cdot \frac{\cos^2(\theta)}{\sin^2(\theta)}\]

OpenStudy (anonymous):

No this is the problem: sin^2 Θcsc^2 Θ = sin^2 Θ = cos^2 Oh well, I am going to ask my teacher if she wrote the equation wrong

OpenStudy (mertsj):

\[\sin ^2\theta \times\frac{1}{\sin ^2\theta}- \sin ^2\theta=\cos ^2\theta\]

OpenStudy (mertsj):

She did.

myininaya (myininaya):

she totally wrote it wrong so i totally agree with mr. mert here she meant to write the way we said just change it you might be only one who does the problem (And might get bonus :))

OpenStudy (anonymous):

Okay I will do that Thank You:)